The sum of julio's and Ramona's ages is 36.twice Julio's age is Ramona's age. Find how old they are
Answer by solve_for_x(190)

Let J represent Julio's age, and let R represent Ramona's age.

Since the sum is 36, you can write:

J + R = 36

Then, since Julio's age is twice Ramona's age, you can write:

J = 2R

Substituting 2R in place of J in the first equation gives:

2R + R = 36

3R = 36

R = 12

Then J = 2R = 2(12) = 24

Solution: Julio's age is 24, and Ramona's age is 12.